# Enhancing Graph-based Recommendations with Majority-Voting LLM-Rerank Augmentation

### v1

@@ Abstract
Recommendation systems often suffer from data sparsity caused by limited user-item interactions, which degrade their performance and amplify popularity bias in real-world scenarios. This paper proposes a novel data augmentation framework that leverages Large Language Models (LLMs) and item textual descriptions to enrich interaction data. By few-shot prompting LLMs multiple times to rerank items and aggregating the results via majority voting, we generate high-confidence synthetic user-item interactions, supported by theoretical guarantees based on the concentration of measure. To effectively leverage the augmented data in the context of a graph recommendation system, we integrate it into a graph contrastive learning framework to mitigate distributional shift and alleviate popularity bias. Extensive experiments show that our method improves accuracy and reduces popularity bias, outperforming strong baselines.
@@ 1 Introduction
Recommendation systems (RS), particularly graph-based recommendation models, often suffer from data sparsity, which significantly degrades performance and amplifies popularity bias (Idrissi and Zellou 2020). Traditional data augmentation techniques such as random walk sampling, node (edge) dropout (Wu et al. 2021), and subgraph sampling (Fan et al. 2023a; Li et al. 2023) generate alternative views of the interaction graph for use in self-supervised learning. These methods are simple and can improve accuracy in low-interaction regimes, but they rely solely on interaction data and overlook the potential of other modalities, such as item textual descriptions. Recent advances leverage the reasoning capabilities of Large Language Models (LLMs) and the rich semantic information in item descriptions to enhance user and item profiling for data augmentation (Wu et al. 2024b).
@@
While LLM-based methods show promise in enriching recommendation datasets using external knowledge, they suffer from two main limitations. First, the augmentation results generated by prompting LLMs are often inconsistent across different inference runs, leading to unstable performance. (Hou et al. 2024b; Wei et al. 2024). Specifically, using a commonly adopted prompt template from prior research to augment interaction data, the results on the Netflix dataset sampled with 100010001000 users to reduce augmentation cost show that LLMRec (Hou et al. 2024b) and Llama4Rec (Luo et al. 2024) yield inconsistent NDCG@10 scores. This metric evaluates how well the ranking of the top 10 items reflects user preferences, as illustrated in Figure 1.
Figure 1: NDCG@10 changes over LLM augmentation runs
Figure 2: NDCG@10 of LLM-based reranking over multiple inference runs. Red dashed line: LightGCN baseline; green solid line: ensemble result of LLM runs on 2 datasets.
This is because the augmented items vary across different runs, leading to unstable outcomes. Second, embeddings derived from the textual content of LLM-embedded models tend to be out-of-distribution compared to collaborative embeddings, making it difficult to align them in a model-agnostic framework across various RS (Ren et al. 2024; Qiao et al. 2024).
To address these issues, we propose a reformulated prompting strategy that casts the augmentation task as an item reranking problem, enhanced with a few-shot prompting design. This format enables the LLM to better leverage its parametric knowledge when ranking candidate items according to user preferences, encouraging the model to assign high positive scores to items that are more likely to match user interests (Wu et al. 2024a; Hou et al. 2024c). As shown in Figure 2, although individual LLM reranking runs yield varying NDCG@10 scores, they outperform the traditional graph-based method LightGCN (He et al. 2020) in most cases. Aggregating multiple reranking outputs via majority voting further improves and stabilizes performance, demonstrating the LLM’s capacity to adapt to user preferences. To quantify this effect, we vary the number of voting runs and conduct 30 repeated inferences for each setting.
Figure 3: Distribution of NDCG@10 over different number of ensemble
The results, visualized in Figure 3, show that increased voting numbers improves mean NDCG@10 and reduces variance, indicating greater robustness.
Motivated by this, we propose a theoretically grounded data augmentation strategy based on majority-vote LLM reranking, supported by concentration of measure theory to bound augmentation accuracy. Instead of incorporating LLM-generated textual embeddings, we directly leverage the augmented data via a contrastive learning (CL) objective applied between original and augmented graphs. This enables representation alignment while mitigating distributional shift and popularity bias, without modifying the base architecture. We summarize our main contributions as follows: (1), we propose VoteGCL, a model-agnostic, end-to-end framework that employs few-shot prompted LLMs for in-context reranking. By performing majority-vote reranking over candidate items, VoteGCL identifies high-confidence user-item interactions for data augmentation. It further integrates graph contrastive learning (CL) between original and augmented views to exploit the augmented data and address distributional shift; (2) we provide a theoretical analysis based on the concentration of measure, demonstrating that majority-vote reranking effectively extracts the LLM’s most reliable knowledge, thereby enhancing the quality of augmented interactions; and (3) extensive experiments show that VoteGCL consistently outperforms state-of-the-art graph-based, multi-modal, and LLM-enhanced recommender models across multiple benchmarks, improving recommendation accuracy while mitigating popularity bias.
@@ 2 Related work
Graph-Augmented Recommendation. Graph-based RS represent user-item interactions as bipartite graphs, where nodes are users and items, and edges denote interactions. Early models used random walks to uncover latent relationships via transition probabilities (Nikolakopoulos and Karypis 2019). The introduction of Graph Convolutional Networks (GCNs) enabled embedding propagation to better capture collaborative signals (He et al. 2020; Wang et al. 2019). Various graph augmentation techniques have been developed to address data sparsity and over-smoothing issues in graph neural networks. Traditional graph augmentation methods (Ding et al. 2022) can be categorized into three approaches: (1) Edge/Node Dropout, which removes redundant connections to improve robustness (Wu et al. 2021; Zhou et al. 2023); (2) Graph Diffusion, which enriches connectivity by adding new edges based on structural properties (Fan et al. 2023b; Yang et al. 2021); and (3) Subgraph Sampling, which generates localized views for diverse training samples (Zhang et al. 2024; Xie et al. 2022). While effective, these methods focus solely on graph structure and overlook valuable side information such as item content, motivating more holistic augmentation strategies.
@@ 3 Preliminaries
Problem Formulation. In graph recommendation, user-item interactions are modeled as a bipartite graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$ includes user nodes $\mathcal{U}$ and item nodes $\mathcal{I}$, and $\mathcal{E}$ denotes observed interactions. Each user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$ is associated with a learnable embedding $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d}$, initialized in the embedding table $\mathbf{E}^{(0)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ and updated during training, where $d$ is the embedding size. The training process is performed on mini-batches $\mathcal{B}$ of user-item interactions sampled from $\mathcal{E}$.
Vanilla Graph-based Retrieval. We adopt LightGCN (He et al. 2020), a simplified yet effective GCN, as the primary retrieval model to propagate information over the user-item graph $\mathcal{G}$ by updating node representations through neighborhood aggregation and combination. At the $l$-th iteration, the node representations are updated as: $\mathbf{E}^{(l)}=H(\mathbf{E}^{(l-1)},\mathcal{G}),$ where $\mathbf{E}^{(l)}$ denotes the node representations at iteration $l$ which corresponds to the $l$-th layer of graph convolution, and $H$ is the neighborhood aggregation function. The embedding propagation process is mathematically formulated as a weighted combination of representations from different iterations. Specifically, the final node embeddings are computed as:
@@
where $\mathbf{E}$ denotes the final representations used for downstream prediction tasks, $L$ corresponds to the number of propagation layers, and $\tilde{\mathbf{A}}\in\mathbf{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the normalized undirected adjacency matrix without self-connections. This iterative process allows the model to capture higher-order connectivity patterns in the graph. A common approach is to train the graph-based recommendation model using the Bayesian Personalized Ranking (BPR) loss. The model learns to assign higher scores to observed interactions than to unobserved ones. The predicted score for user $u$ and item $i$ is computed via inner product: $\hat{y}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}.$ The BPR loss is defined as:
@@
where $O=\left\{(u,i,j)\mid(u,i)\in\mathcal{E},(u,j)\in\mathcal{R}^{-}\right\}$ denotes the pairwise training data, $\mathcal{R}^{-}$is the unobserved interactions and $\sigma(\cdot)$ is the sigmoid function. During training, the user embedding $\mathbf{e}_{u}$ is optimized to be close to the embeddings $\mathbf{e}_{i}$ of preferred items. In inference, recommendations are made by selecting items whose embeddings are nearest to $\mathbf{e}_{u}$.
@@ 4 Methodology
We propose framework VoteGCL, as illustrated in Figure 4, we use few-shot prompting LLM for inference via majority-voted reranking to construct the augmented dataset as detailed in Section 4.1. To effectively leverage the augmented data for graph recommendation, we employ the CL paradigm as detailed in Section 4.2.
4.1 Data Augmentation via Majority LLM Reranking
Augmenting artificial interactions for high-degree users can introduce unnecessary cost and noisy interactions. Therefore, we apply the augmentation only to low-degree users, defined as those whose number of interactions falls below the 25th percentile ($Q_{\alpha}=0.25$) of the user interaction distribution. This strategy ensures that edges are augmented only for users with insufficient interaction data. An ablation study analyzing the effect of this threshold is provided in Section 5.3.
For each user $u$ in the selected set, we aim to select $p$ new interactions that do not appear in the original dataset. On the other hand, while LLMs have demonstrated greater effectiveness, they are limited by context length and thus cannot consider all potential items and related users within a single prompt. Hence, we first propose applying a conventional retrieval model to obtain an ordered shortlist of $K$ non-interacted candidate items, denoted as $\textbf{I}_{\text{can}}=\{i_{k}\}_{k=1}^{K}$, along with a set of similar users based on collaborative embeddings generated by the retrieval model, inspired by Wu et al.; Wei et al. Based on this information, we employ an LLM to rerank the sequence $\textbf{I}_{\text{can}}$, producing a sorted list denoted as $\textbf{I}_{\text{sorted}}$. To perform the reranking, we design a generalizable few-shot prompting template consisting of four components: first, a task-specific instruction; second, the user’s watch history enriched with flattened metadata; third, a structured output format; and fourth, one or more few-shot examples drawn from similar users. Each example includes the user’s history along with a ground-truth reranked list sorted by rating. This prompt is dataset-agnostic and can be easily adapted to new domains by reformatting the metadata. An example prompt for the movie dataset is provided in Figure 4, with further details available in Appendix LABEL:prompt.
To improve the consistency of reranking outcomes and better leverage the LLM’s reasoning capabilities across multiple trials, we perform $N$ independent rerankings of the candidate list $\textbf{I}_{\text{can}}$. The resulting sequences are then aggregated using a simplified variant of Reciprocal Rank Fusion (Chapelle et al. 2009). For each item $i_{k}\in\textbf{I}_{\text{can}}$, we compute an aggregated score as follows:
@@ 4.1 Data Augmentation via Majority LLM Reranking
where $\varsigma^{(n)}(i_{k})$ denotes the position of item $i_{k}$ in the $n$-th reranked sequence. We select the top $p$ items with the highest scores to construct new artificial interactions for the user $u$. We repeat the process for all targeted users, leading to $Q_{\alpha}\times|\mathcal{V}|\times p$ newly augmented edges. After data augmentation, we obtain an enriched graph $\mathcal{G}_{\text{aug}}=(\mathcal{V},\mathcal{E}^{+})$, where the original edge set $\mathcal{E}$ is augmented with newly generated interactions, thereby expanding $\mathcal{E}^{+}=\mathcal{E}\cup\mathcal{E}_{\text{new}}$ while preserving the bipartite structure between users $\mathcal{U}$ and items $\mathcal{I}$.
@@
Let =$\mathcal{I}=\{i_{1},i_{2},\ldots,i_{K}\}$ be a set of $K$ items, and let $\mathcal{T}_{K}$ denote all permutations of $\{1,2,\ldots,K\}$. Let $P$ be any probability distribution over $\mathcal{T}_{K}$, and let $\varsigma^{(1)},\ldots,\varsigma^{(N)}$ be independent random permutations sampled from $P$. Let $g:\mathbb{R}\to\mathbb{R}$ be a bounded, strictly decreasing function. Suppose $A,B\in\mathbb{R}$ are such that $A\leq g(x)\leq B$ for all $x$. For each item $i\in\mathcal{I}$, define the aggregated score
$S(i)=\sum_{n=1}^{N}g\big{(}\varsigma^{(n)}(i)\big{)}.$
Fix distinct items $i_{j},i_{k}\in\mathcal{I}$, and let
$\mu=\mathbb{E}_{\varsigma\sim P}\left[g\big{(}\varsigma(i_{j})\big{)}-g\big{(}\varsigma(i_{k})\big{)}\right].$
If $\mu<0$, then
@@
This theorem is sufficent for $g(x)$ in Equation 3 and $\forall i_{j},i_{k}\in\mathcal{I}$, let $\mu=\mathbb{E}\left[\frac{1}{\varsigma(i_{j})+1}-\frac{1}{\varsigma(i_{k})+1}\right].$ If $\mu<0$, then
@@
Theorem 1 illustrates that with a sufficiently large number of LLM reranking steps $N$, the ensemble output converges toward the LLM’s optimal parametric knowledge for the reranking task, thereby yielding high-confidence augmented data. We provided detailed proof of Theorem 1 in Appendix LABEL:theorem.
@@
Unlike other LLM-augmented data approaches summarized in Table 1, VoteGCL does not incur the cost of generating textual embeddings for user/item profiling or adapting high-dimensional textual representations into recommendation models. Instead, VoteGCL focuses on generating high-quality interaction data and effectively leveraging it through a Graph CL paradigm, as detailed in Section 4.2. The key advantages of our LLM-augmented data are twofold: (1) a majority-vote reranking mechanism, which ensures more consistent and high-confidence user-item interactions, with detail theoretical justification provided in Appendix LABEL:theorem; and (2) a few-shot prompting strategy that incorporates collaborative signals from similar users, enabling better adaptation to user-specific preferences in the data augmentation task.

### v3

@@ Abstract
Recommendation systems often suffer from data sparsity, caused by limited user-item interactions, which degrades their performance and amplifies popularity bias in real-world scenarios. This paper proposes a novel data augmentation framework that leverages Large Language Models (LLMs) and item textual descriptions to enrich interaction data. By few-shot prompting LLMs multiple times to rerank items and aggregating the results via majority voting, we generate high-confidence synthetic user-item interactions, supported by theoretical guarantees based on the concentration of measure. To effectively leverage the augmented data in the context of a graph recommendation system, we integrate it into a graph contrastive learning framework to mitigate distributional shift and alleviate popularity bias. Extensive experiments show that our method improves accuracy and reduces popularity bias, outperforming strong baselines.
@@ 1 Introduction
Recommendation systems (RS), particularly graph-based recommendation models, often suffer from data sparsity, which significantly degrades performance and amplifies popularity bias (Idrissi and Zellou 2020). Traditional data augmentation techniques such as random walk sampling, node and edge dropout (Wu et al. 2021), and subgraph sampling (Fan et al. 2023a; Li et al. 2023) generate alternative views of the interaction graph for use in self-supervised learning. These methods are simple and can improve accuracy in low-interaction regimes. Still, they rely solely on interaction data and overlook the potential of other modalities, such as item textual descriptions.
Figure 1: NDCG@10 variation across repeated LLM augmentation runs on 1,000 Netflix users using a casual prompt. Variability in augmented items across runs leads to unstable recommendation quality. Higher NDCG@10 indicates better alignment with user preferences.
Figure 2: Distribution of NDCG@10 across different reranking times over 30 repeated runs. Increasing the ensemble size via more reranking times improves the mean and reduces variance, indicating greater stability and robustness. LLM-based reranking consistently outperforms LightGCN, and ensemble methods further improve accuracy.
@@
Recent advances leverage the reasoning capabilities of Large Language Models (LLMs) and the rich semantic information in item descriptions to enhance user and item profiling for data augmentation (Wu et al. 2024b). While LLM-based methods show promise in enriching recommendation datasets using external knowledge, they suffer from two main limitations. First, the augmentation results generated by prompting LLMs are often inconsistent across different inference runs, leading to unstable performance. Specifically, using a commonly adopted prompt template from prior works to augment interaction data, we observe that both LLMRec (Hou et al. 2024b) and Llama4Rec (Luo et al. 2024) yield fluctuating NDCG@10 scores on the Netflix dataset, as shown in Figure 1. Second, embeddings derived from the textual content of LLM-embedded models tend to be out-of-distribution compared to collaborative embeddings, making it difficult to align them in a model-agnostic framework across various RS (Ren et al. 2024; Qiao et al. 2024).
To address the above challenges, we reformulate LLM-based data augmentation as an item reranking task, combining few-shot prompting with majority-vote ensembling via Reciprocal Rank Fusion (RRF) (Cormack, Clarke, and Buettcher 2009). This formulation enables the LLM to better leverage its parametric knowledge of user preferences, assigning high positive scores to relevant items (Wu et al. 2024a; Hou et al. 2024c). Ensemble inference improves robustness and consistency by aggregating signals across reranking runs, as shown in Figures  2. To support this ensembling approach, we propose a theoretically grounded augmentation strategy, using concentration of measure theory to derive bounds on augmentation accuracy. Rather than relying on LLM-generated embeddings, we directly utilize the augmented data in a graph contrastive learning (CL) framework. This approach aligns representations between original and augmented graphs, mitigating distributional shift and popularity bias (Zimmermann et al. 2021), without modifying the base model architecture. We summarize our contributions as:
  • We propose VoteGCL, a model-agnostic, end-to-end framework that employs LLMs with few-shot in-context reranking. VoteGCL uses majority-vote reranking over candidate items to perform data augmentation and further integrates graph contrastive learning to effectively leverage the augmented data.
  • We provide a theoretical analysis based on the concentration of measure, demonstrating that majority-vote reranking effectively extracts the LLM’s most reliable knowledge, thereby enhancing the quality of augmented interactions.
  • Extensive experiments show that VoteGCL consistently outperforms state-of-the-art graph-based and LLM-enhanced recommender models across multiple benchmarks, improving recommendation accuracy while mitigating popularity bias.
@@ 2 Related work
Graph-Augmented Recommendation. Graph-based RS represents user-item interactions as bipartite graphs, where nodes are users and items, and edges denote interactions. Early models used random walks to uncover latent relationships via transition probabilities (Nikolakopoulos and Karypis 2019). The introduction of Graph Convolutional Networks (GCNs) enabled embedding propagation better to capture collaborative signals (He et al. 2020; Wang et al. 2019). Various graph augmentation techniques have been developed to address data sparsity and over-smoothing issues in graph RS. Traditional graph augmentation methods (Ding et al. 2022) can be categorized into three approaches: (1) Edge/Node Dropout, which removes redundant connections to improve robustness (Wu et al. 2021; Zhou et al. 2023); (2) Graph Diffusion, which enriches connectivity by adding new edges based on structural properties (Fan et al. 2023b; Yang et al. 2021); and (3) Subgraph Sampling, which generates localized views for diverse training samples (Zhang et al. 2024; Xie et al. 2022). While effective, these methods focus solely on graph structure and overlook valuable side information such as item content.
@@ 3 Preliminaries
Problem Formulation. In graph-based recommendation systems, user-item interactions are modeled as a bipartite graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\mathcal{U}\cup\mathcal{I}$ consists of user nodes $\mathcal{U}$ and item nodes $\mathcal{I}$, and $\mathcal{E}$ denotes the set of observed interactions. Each user $u\in\mathcal{U}$ and item $i\in\mathcal{I}$ is assigned a learnable embedding $\mathbf{e}_{u},\mathbf{e}_{i}\in\mathbb{R}^{d}$, initialized in an embedding matrix $\mathbf{E}^{(0)}\in\mathbb{R}^{|\mathcal{V}|\times d}$ and updated during training. The model is optimized on mini-batches $\mathcal{B}$ sampled from $\mathcal{E}$. After training, we denote the resulting embedding matrix as $\mathbf{E}$, which is learned such that each user embedding $\mathbf{e}_{u}$ is close to the embeddings $\mathbf{e}_{i}$ of the items the user prefers.
Vanilla Graph-based Retrieval. We adopt LightGCN (He et al. 2020), a simplified yet effective GCN, as the primary retrieval model to propagate information over the user-item graph $\mathcal{G}$ by updating node representations through neighborhood aggregation and combination. At the $l$-th iteration, the node representations are updated as: $\mathbf{E}^{(l)}=H(\mathbf{E}^{(l-1)},\mathcal{G}),$ where $\mathbf{E}^{(l)}$ denotes the node representations at iteration $l$ which corresponds to the $l$-th layer of graph convolution, and $H$ is the neighborhood aggregation function. The embedding propagation process is mathematically formulated as a weighted combination of representations from different iterations. The final embeddings are computed as:
@@
where $\mathbf{E}$ denotes the final representations used for downstream prediction tasks, $L$ corresponds to the number of propagation layers, and $\tilde{\mathbf{A}}\in\mathbf{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ is the normalized undirected adjacency matrix without self-connections. This iterative process allows the model to capture higher-order connectivity patterns in the graph. A common approach is to train the graph-based recommendation model using the Bayesian Personalized Ranking (BPR) (Rendle et al. 2012) loss. The model learns to assign higher scores to observed interactions than to unobserved ones. The predicted score for user $u$ and item $i$ is computed via inner product: $\hat{y}_{ui}=\mathbf{e}_{u}^{\top}\mathbf{e}_{i}.$ The BPR loss is defined as:
@@
where $O=\left\{(u,i,j)\mid(u,i)\in\mathcal{E},(u,j)\in\mathcal{R}^{-}\right\}$ denotes the pairwise training data, $\mathcal{R}^{-}$is the unobserved interactions and $\sigma(\cdot)$ is the sigmoid function. In inference, items closest to $\mathbf{e}_{u}$ in embedding space are recommended.
@@ 4 Methodology
We propose a framework VoteGCL, as illustrated in Figure 3, we use few-shot prompting LLM for inference via majority-voted reranking to construct the augmented dataset as detailed in Section 4.1. To effectively leverage the augmented data for graph recommendation, we employ the CL paradigm as detailed in Section 4.2.
4.1 Data Augmentation via Majority Reranking
This section presents a targeted data augmentation strategy aimed at enriching the training graph with high-quality synthetic interactions. Augmenting artificial interactions for high-degree users can introduce unnecessary cost and noisy interactions. Therefore, we apply the augmentation only to low-degree users, defined as those whose number of interactions falls below the 25th percentile of the user interaction distribution. This strategy ensures that edges are augmented only for users with insufficient interaction data. An ablation study analyzing the effect of this threshold is provided in Section 5.3.
For each user $u$ in the selected set, we aim to select $p$ new interactions that do not appear in the original dataset. While LLMs have shown promising capabilities in ranking and recommendation tasks, they are limited by context length and thus cannot consider all potential items and related users within a single prompt. Hence, we employ LightGCN as a conventional retrieval model to obtain an ordered shortlist of $K$ non-interacted candidate items, denoted as $\textbf{I}_{\text{can}}=\{i_{k}\}_{k=1}^{K}$, along with a set of similar users based on collaborative embeddings generated by the retrieval model, inspired by Wu et al.; Wei et al.. Based on this information, we employ an LLM to rerank the sequence $\textbf{I}_{\text{can}}$, producing a sorted list denoted as $\textbf{I}_{\text{sorted}}$. To perform the reranking, we design a generalizable few-shot prompting template consisting of four components: first, a task-specific instruction; second, the user’s watch history enriched with flattened metadata; third, a structured output format; and fourth, one or more few-shot examples drawn from similar users. Each example includes the user’s history along with a ground-truth, reranked list sorted by rating. This prompt is dataset-agnostic and can be easily adapted to new domains by reformatting the metadata. An example prompt for the movie dataset is provided in Figure 3, with further details available in Appendix C.
To improve the consistency of reranking outcomes and better leverage the LLM’s reasoning capabilities across multiple trials, we perform $N$ independent rerankings of the candidate list $\textbf{I}_{\text{can}}$. The resulting sequences are then aggregated using a simplified variant of RRF. For each item $i_{k}\in\textbf{I}_{\text{can}}$, we compute an aggregated score as follows:
@@ 4.1 Data Augmentation via Majority Reranking
where $\varsigma^{(n)}(i_{k})$ denotes the position of item $i_{k}$ in the $n$-th reranked sequence. We select the top $p$ items with the highest scores to construct new artificial interactions for the user $u$. Following this data augmentation strategy, we theoretically prove that for any bounded, strictly decreasing function $g(\cdot)$, increasing the number of independent re-rankings $N$ asymptotically reveals the LLM’s optimal parametric knowledge for the reranking task. The formal statement is provided in Theorem 1, and a detailed proof can be found in Appendix D.
@@
Let $\mathcal{I}=\{i_{1},i_{2},\ldots,i_{K}\}$ be a set of $K$ items, and let $\mathcal{T}_{K}$ denote all permutations of $\{1,2,\ldots,K\}$. Let $P$ be any probability distribution over $\mathcal{T}_{K}$, and let $\varsigma^{(1)},\ldots,\varsigma^{(N)}$ be independent random permutations sampled from $P$. Let $g:\mathbb{R}\to\mathbb{R}$ be a bounded, strictly decreasing function. Suppose $A,B\in\mathbb{R}$ are such that $A\leq g(x)\leq B$ for all $x$. For each item $i\in\mathcal{I}$, define the aggregated score
$S(i)=\sum_{n=1}^{N}g\big{(}\varsigma^{(n)}(i)\big{)},$
where $\varsigma^{(n)}(i)$ is the rank assigned to item $i$ by permutation $\varsigma^{(n)}$. Let consider two distinct items $i_{j},i_{k}\in\mathcal{I}$, and $\mu=\mathbb{E}_{\varsigma\sim P}\left[\varsigma(i_{j})-\varsigma(i_{k})\right]$. If $\mu>0$, then
@@
We instantiate $g(\cdot)$ as in Equation 3 for its simplicity and prominence in prior work. More importantly, given a finite value of $K$, the $g(\cdot)$ function is strictly decreasing and bounded, satisfying the requirements of Theorem 1. The corollary below specializes in the result for this choice, and its proof is presented in Appendix D.
Corollary 1 (Reciprocal Rank Aggregation).
Let $g(x)=1/(x+1)$ and define $S(i)=\sum_{n=1}^{N}1/(\varsigma^{(n)}(i)+1)$ for each $i\in\mathcal{I}$. For any $i_{j},i_{k}\in\mathcal{I}$, let $\mu=\mathbb{E}_{\varsigma\sim P}\left[\varsigma(i_{j})-\varsigma(i_{k})\right]$. If $\mu>0$, then
@@
The probability distribution $P$ appearing in both the Theorem and Corollary is genuinely the LLM-induced distribution over permutations. Theorem 1 and Corollary 1 show that when item $i_{k}$ is expected to be ranked higher than item $i_{j}$, meaning it has a lower average rank and $\mu>0$, the probability that $i_{j}$ receives a higher aggregated score than $i_{k}$ decreases exponentially with the number of permutations $N$. As $N\to\infty$, this error probability vanishes, indicating that score aggregation becomes increasingly reliable. Therefore, selecting the top $p$ items with the highest aggregated scores after enough rerankings recovers the optimal set of $p$ items under the LLM’s ideal parametric ranking.
We repeat the process for all targeted users, leading to $Q_{\alpha}\times|\mathcal{V}|\times p$ newly augmented edges. After data augmentation, we obtain an enriched graph $\mathcal{G}_{\text{aug}}=(\mathcal{V},\mathcal{E}^{+})$, where the original edge set $\mathcal{E}$ is augmented with newly generated interactions, thereby expanding $\mathcal{E}^{+}=\mathcal{E}\cup\mathcal{E}_{\text{new}}$ while preserving the bipartite structure between users $\mathcal{U}$ and items $\mathcal{I}$.
@@
Unlike other LLM-augmented data approaches summarized in Table 1, VoteGCL does not incur the cost of generating textual embeddings for user/item profiling or adapting high-dimensional textual representations into recommendation models. Instead, VoteGCL focuses on generating high-quality interaction data and effectively leveraging it through a Graph CL paradigm, as detailed in Section 4.2. The key advantages of our LLM-augmented data are twofold: (1) a majority-vote reranking mechanism, which ensures more consistent and high-confidence user-item interactions, with detailed theoretical justification provided in Appendix D; and (2) a few-shot prompting strategy that incorporates collaborative signals from similar users, enabling better adaptation to user-specific preferences in the data augmentation task.